OpenStudy (anonymous):

displacement time equation of a particle moving along x-axis is x=20+t3-12t. a)find velocity of particle at time t=0 b)state whether the motion is uniformly acceleration or nt. c)find position of the particle when v=0.

OpenStudy (apoorvk):

if that^ is the equation of the displacement, how would you get the equation for velocity?

OpenStudy (apoorvk):

wrt 'time' ofcourse.

OpenStudy (anonymous):

"Substitute" the value of "t" and find.

OpenStudy (anonymous):

Hint for part (a): \(v=\frac{dx}{dt}\) Hint for part (b): \(a=\frac{dv}{dt} = \frac{d^2x}{dt^2}\). If this is a function of time, then it is not constant. Hint for part (c): What \(t\) values satisfies this condition?

OpenStudy (anonymous):

v=dx/dt

OpenStudy (anonymous):

Sorry. @yakeyglee is absolutely correct. Follow his steps and ur answer is done

OpenStudy (anonymous):

as v=0,we 0=3t2-12

OpenStudy (anonymous):

am i right.

OpenStudy (anonymous):

Yeah, perfect @Ruchi.

OpenStudy (anonymous):

So, we get